OpenStudy (anonymous):

Which of the following is equivalent to |x + 11| = 13? x + 11 = 13 x + 11 = - 13 x + 11 = 13 or x + 11 = -13 - x + 11 = 13

OpenStudy (paxpolaris):

\(x + 11 = 13\) or \(x + 11 = -13\)

OpenStudy (paxpolaris):

since... \(\large \left| 13 \right|= 13\) and \(\large \left| -13 \right|= 13\)

OpenStudy (jiteshmeghwal9):

\[X+11=13\] \[x=13-11\] \[x=2\]